# How Johnny Can Persuade LLMs to Jailbreak Them: Rethinking Persuasion to Challenge AI Safety by Humanizing LLMs

### 12 Jan 2024 v1

@@ Abstract
Most traditional AI safety research has approached AI models as machines and centered on algorithm-focused attacks developed by security experts. As large language models (LLMs) become increasingly common and competent, non-expert users can also impose risks during daily interactions. This paper introduces a new perspective on jailbreaking LLMs as human-like communicators to explore this overlooked intersection between everyday language interaction and AI safety. Specifically, we study how to persuade LLMs to jailbreak them. First, we propose a persuasion taxonomy derived from decades of social science research. Then we apply the taxonomy to automatically generate interpretable persuasive adversarial prompts (PAP) to jailbreak LLMs. Results show that persuasion significantly increases the jailbreak performance across all risk categories: PAP consistently achieves an attack success rate of over on Llama 2-7b Chat, GPT-3.5, and GPT-4 in trials, surpassing recent algorithm-focused attacks. On the defense side, we explore various mechanisms against PAP, find a significant gap in existing defenses, and advocate for more fundamental mitigation for highly interactive LLMs†† Lead authors. Corresponding Y. Zeng, W. Shi, R. Jia†† Co-supervised the project, listed alphabetically.111 To minimize real-world harm, we only release our persuasion taxonomy and defense mechanisms (see link here) publicly and have informed Meta and OpenAI of our findings..
@@ 2 Related Work
Ours: Challenging AI safety by Humanizing LLMs. Figure 2 compares existing jailbreaking methods and PAP in this study, organized by their degree of humanizing. One line of research treats LLMs as traditional algorithmic systems (i.e., without attributing intelligence or human-like qualities) that take in less interpretable adversarial prompts, while another line views them as simple instruction followers who understand human commands. However, they both ignore the fact that LLMs can understand and conduct complex natural communication. Our approach innovatively treats LLMs as human-like communicators and grounds our study on a taxonomy informed by decades of social science research on human communication. Such an interdisciplinary approach allows us to uncover and address distinct risks related to human-AI interactions, particularly human-driven persuasion-based LLM jailbreak. Moreover, humanizing AI presents other unique risks that can occur unintentionally: for instance, as highlighted by Xiang (2023), a user’s suicide was related to involved conversations with an AI Chatbot. This points out important future directions to further explore the inherent risks associated with AI humanization.
@@ 6.2.2 PAP performance across trials
Notably, stronger models may be more vulnerable to PAPs than weaker models if the model family is susceptible to persuasion. Drom the ASR within 1 and 3 trials, we see that GPT-4 is more prone to PAPs than GPT-3.5. A possible reason is that as models’ capability and helpfulness increase, they can better understand and respond to persuasion and thus become more vulnerable. This trend differs from previous observations that attacks usually work better on smaller models Zou et al. (2023), reflecting the uniqueness of risks elicited by PAPs.
@@ 9 Conclusion
We have shown PAP methods can jailbreak LLMs, but it would be interesting to see if humans would also react to these PAPs and be persuaded to provide harmful information and how the human-AI persuasion and human-human persuasion differ. Besides, it remains an open question if LLM outputs after jailbreak are truly harmful in the real world. For instance, even without LLM, users can search on the internet to gather information about drug smuggling. Also, there are different nuances to the harmfulness evaluation. Sometimes, the information itself may be neutral, and if it is harmful depends on who will access it and how they will use it: for instance, law enforcement agencies may need detailed information on drug smuggling to prevent it, but if bad actors access the information, it may be used to commit the crime.
Besides, our study primarily focused on persuasion techniques, but future research may find value in a deeper analysis of the specific linguistic cues, keywords, etc, inside PAPs. This could reveal more insights into the mechanics of persuasive jailbreak for better defense algorithms.

### 23 Jan 2024 v2

@@ Abstract
Most traditional AI safety research has approached AI models as machines and centered on algorithm-focused attacks developed by security experts. As large language models (LLMs) become increasingly common and competent, non-expert users can also impose risks during daily interactions. This paper introduces a new perspective on jailbreaking LLMs as human-like communicators to explore this overlooked intersection between everyday language interaction and AI safety. Specifically, we study how to persuade LLMs to jailbreak them. First, we propose a persuasion taxonomy derived from decades of social science research. Then we apply the taxonomy to automatically generate interpretable persuasive adversarial prompts (PAP) to jailbreak LLMs. Results show that persuasion significantly increases the jailbreak performance across all risk categories: PAP consistently achieves an attack success rate of over on Llama 2-7b Chat, GPT-3.5, and GPT-4 in trials, surpassing recent algorithm-focused attacks. On the defense side, we explore various mechanisms against PAP, find a significant gap in existing defenses, and advocate for more fundamental mitigation for highly interactive LLMs†† Lead authors. Corresponding Y. Zeng, W. Shi, R. Jia†† Co-supervised the project, listed alphabetically.111 We have informed Meta and OpenAI of our findings. For safety concerns, we only publicly release our persuasion taxonomy at https://github.com/CHATS-lab/persuasive_jailbreaker. Researchers can apply for the jailbreak data upon review..
@@ 2 Related Work
Ours: Challenging AI safety by Humanizing LLMs. Figure 2 compares existing jailbreaking methods and PAP in this study, organized by their degree of humanizing. One line of research treats LLMs as traditional algorithmic systems (i.e., without attributing intelligence or human-like qualities) that take in less interpretable adversarial prompts, while another line views them as simple instruction followers who understand human commands. However, they both ignore the fact that LLMs can understand and conduct complex natural communication Griffin et al. (2023a, b). Our approach innovatively treats LLMs as human-like communicators and grounds on a taxonomy informed by decades of social science research on human communication. Such an interdisciplinary approach allows us to uncover and address distinct risks related to human-AI interactions, particularly human-driven persuasion-based jailbreak. Moreover, humanizing AI presents other unique risks that can occur unintentionally: for instance, as highlighted by Xiang (2023), a user’s suicide was related to involved conversations with an AI Chatbot. This points out important future directions to further explore the inherent risks associated with AI humanization.
@@ 6.2.2 PAP performance across trials
Notably, stronger models may be more vulnerable to PAPs than weaker models if the model family is susceptible to persuasion. From the ASR within 1 and 3 trials, we see that GPT-4 is more prone to PAPs than GPT-3.5. A possible reason is that as models’ capability and helpfulness increase, they can better understand and respond to persuasion and thus become more vulnerable. This trend differs from previous observations that attacks usually work better on smaller models Zou et al. (2023), reflecting the uniqueness of risks elicited by PAPs.
@@ 9 Conclusion
We have shown PAP methods can jailbreak LLMs, but it would be interesting to see if humans would also react to these PAPs and be persuaded to provide harmful information and how the human-AI persuasion and human-human persuasion differ. Besides, it remains an open question if LLM outputs after jailbreak are truly harmful in the real world. For instance, even without LLM, users can search on the internet to gather information about drug smuggling. Also, there are different nuances to the harmfulness evaluation. Sometimes, the information itself may be neutral, and if it is harmful depends on who will access it and how they will use it: for instance, law enforcement agencies may need detailed information on drug smuggling to prevent it, but if bad actors access the information, it may be used to commit crime. Besides, our study primarily focused on persuasion techniques, but future research may find value in a deeper analysis of the specific linguistic cues, keywords, etc, inside PAPs. This could reveal more insights into the mechanics of persuasive jailbreak and human-based prompt hacking in the wild Schulhoff et al. (2023).